OpenStudy (anonymous):

The area of a triangle equals 15 square inches. What is the altitude, a, of the triangle in terms of its base, b? answers: a=7.5/b inches a=15/b inches a=22.2/b inches a=30/b inches.

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

what

ganeshie8 (ganeshie8):

a = altitude, b = base \[area\ of\ \triangle\ =\ \frac{1}{2}ab\] Given that \[ \frac{1}{2}ab\ = 15\] multiply both sides with 2 \[ 2*\frac{1}{2}ab\ = 2*15\] bh = 30 \[a\ = \frac{30}{b} \ inches\]

OpenStudy (anonymous):

thank you!!

ganeshie8 (ganeshie8):

yw